You are thinking of purchasing a house. The house costs $350,000. You have $50,000 in cash that you can use as a down payment on the house, but you need to borrow the rest of the purchase price. The bank is offering a 30-year mortgage that requires annual payments and has an interest rate of 7% per year. You can afford to pay only $23,500 per year. The bank agrees to allow you to pay this amount each year, yet still borrow $300,000. At the end of the mortgage (in 30 years), you must make a balloon payment; that is, you must repay the remaining balance on the mortgage. How much will this balloon payment be? Please show work on Excel.

a. What is the PV of payments?

b. What is the Loan PV of payments?

c. What is the Balloon payment?
Answer #1
a) PV of payments = 23500*(1.07^30-1)/(0.07*1.07^30) = $     2,91,612
b) Loan PV of payments $     3,00,000
c) Balloon payment required = (300000-291612)*1.07^30 = $         63,852